Solve 3x^2 + 19x-14=0 using the quadratic formula

Answer by stanbon(75887)

Solve 3x^2 + 19x-14=0 using the quadratic formula
------------------------
x = [-19 +- sqrt(19^2 - 4*3*-14)]/6
x = [-19 +- sqrt(529)]/6
x = [-19 +- 23]/6
x = 2/3 or x = -7
====================
Cheers,
Stan H.